OpenStudy (anonymous):

Solve : 3x-y=-9 for y?

OpenStudy (anonymous):

y = 9 + 3x

OpenStudy (anonymous):

add +y on both sides 3x-y+y=-9+y 3x=-9+y now add 9 on both sides 3x+9=-9+9+y 3x+9=y

OpenStudy (anonymous):

Thanks @sami-21

OpenStudy (anonymous):

yw:)